OpenStudy (mr.math):

Find all integers a,b for which \(a^4+4b^4\) is a prime.

OpenStudy (anonymous):

a=b=1 is the only possible solution.

OpenStudy (mr.math):

Could you prove your claim? :-P

hero (hero):

Yeah FFM

OpenStudy (anonymous):

Yes I definitely can ... why you don't believe me MR.Math? :(

OpenStudy (mr.math):

lol, I do believe you! Where did I say that I don't? :-(

hero (hero):

FFM, stop stalling and post the proof :P

OpenStudy (anonymous):

hehe, okay here it goes, \[ a^4+4b^4= ((a+b)^2+b^2) \times ((a-b)^2+b^2)) \] Now, \( ((a+b)^2+b^2) \gt 1\) (always) so for \( a^4+4b^4 \) to be prime \((a-b)^2+b^2) =1\) and this can only happen when \(a=b=1\) (QED)

OpenStudy (mr.math):

Awesome! I've always believed in you son :-)

OpenStudy (anonymous):

Haha, how old are you man? :D

OpenStudy (mr.math):

Very old.

hero (hero):

FFM, how do you come up with such approaches to proofs?

OpenStudy (anonymous):

Hero, I really wonder that myself ...